OpenStudy (anonymous):

in a book with page numbers from 1 to 100, some pages are torn off. The sum of the numbers on the remaining pages is 4949. How many pages are torn off?

OpenStudy (anonymous):

100 points -

OpenStudy (anonymous):

a complete book shud have \[s_{100}=\frac{100(100+1)}{2}=5050\] right

OpenStudy (shubhamsrg):

sum of 1st hundered nos. = 5050 5050-4949 = 101 Now note that for any page, odd no. comes first on the right side and even no. comes at the left. like page nos will be (1,2) -> 1 on right, and when you turn over, 2 will be on left. then there will be (3,4) -> 3 on right, 4 on left. . . likewise. for 1 page, a+ a+1 = 101 or a=50 but a should be odd, hence wrong ans. for 2 pages, 2a+1 + 2b+1 = 101 a+b = 99/2 -->not possible for 3 pages, a+b+c = 49 a,b,c all are odd this is possible hence 3 can be an answer. infact, 3 should be the answer.

OpenStudy (anonymous):

if i tore page 1 and 100=101 5050-101=4049

OpenStudy (shubhamsrg):

you can not tear 1 and 100

OpenStudy (experimentx):

when you tear one page you remove 2 numbers

OpenStudy (anonymous):

yes thanx for that clearance and 3 is correct answer ,why does @shubhamsrg suppose that 3 has to be correct and no oher answer thereof,

OpenStudy (shubhamsrg):

I dont have a concrete proof for that, i just said said by hit and trial, but I am trying to think of a reason

OpenStudy (asnaseer):

I assume the pages being torn off must be sequencial?

OpenStudy (experimentx):

how many pages were there on book originally? i could be 50 or 52 ... assuming there are two blank pages.

OpenStudy (anonymous):

blank or not its still a page so we count it

OpenStudy (asnaseer):

if that is the case, then the sum of the pages torn off wil be given by:\[\frac{n}{2}(2a+(n-1)d)=101\]where n is the number of pages torn off, 'a' is the first page number, and d=1. you coud then make use of the fact that 101 is prime

OpenStudy (shubhamsrg):

I don't think we only tear off consecutive pages.

OpenStudy (anonymous):

yeah we cud tear numbers that are not consecutive,but it seems like 3 is a unique solution anyway

OpenStudy (anonymous):

on the solution it just says 3 and comment "unique"

OpenStudy (asnaseer):

why couldn't it be one page torn off - with 50 on one side and 51 on the other?

OpenStudy (experimentx):

i think it depends on where 1 stars, 50 and 51 may not be on same page.

OpenStudy (anonymous):

probably cos they are not on the same poage

OpenStudy (anonymous):

how do we verify that

OpenStudy (asnaseer):

there doesn't seem to be enough information given to assert that

OpenStudy (shubhamsrg):

for n pages torn, 2(a_1+a_2+a_3 ...a_n) + n = 101 (a_1+a_2+a_3..a_n) = (101-n)/2 a_1,a_2,a_3..a_n all are odd. n also has to be odd so as to make RHS an integer. Another condition is RHS should be odd integer. n=3,7,11.. satisfy this criteria. for n=7, we have a_1 + a_2 .. a_7 = 47 sum of first 7 odd numbers is 49 i.e. greater that 47 hence anything greater than 3 should not be solution .

OpenStudy (shubhamsrg):

every page should have 2 numbers on either side, hence I;d say we can conclude (49,50) will be one leaf and (50,51) the other

OpenStudy (shubhamsrg):

sorry, (49,50) one and (51,52) other

OpenStudy (asnaseer):

but 49+50+51+52 does not equal 101?

OpenStudy (experimentx):

if 1 starts on front page, then 50 and 51 are not on same page. if 1 starts on back of first page then 50 and 51 are on same page. since the sum is odd, it's good to look for odd number of pages.

OpenStudy (shubhamsrg):

right, it isn't, so (49,50) and (51,52) is not a solution.

OpenStudy (asnaseer):

@experimentX I agree - so why isn't 1 a solution? i.e. the page with 50 on one side and 51 on the other

OpenStudy (anonymous):

bcos 1 starts on front page, then 50 and 51 are not on same page

ganeshie8 (ganeshie8):

yeah otherwise front page wont have an index

OpenStudy (anonymous):

so what pages are this ????

OpenStudy (anonymous):

i mean what pages are torn?

OpenStudy (experimentx):

ok ... since you clarified i suppose it's better to look the solution of the form http://www.wolframalpha.com/input/?i=integer+solutions+%282n%2B1%29%2B%282n%2B2%29%2B%282m%2B1%29%2B%282m%2B2%29%2B%282o%2B1%29%2B%282o%2B2%29+%3D+101

OpenStudy (anonymous):

so there are many possibilities for a combination of 3 pages?

OpenStudy (shubhamsrg):

right, many.

OpenStudy (shubhamsrg):

should be many I mean

OpenStudy (asnaseer):

I see - so we can assert that 3 pages have been torn off, but which 3 cannot be determined from the information given.

OpenStudy (anonymous):

you mean which are obviusly not torn?

OpenStudy (anonymous):

is it can or cannot be dertemined

OpenStudy (experimentx):

also consider if 5 pages might have been torn off, i haven't checked it out.

OpenStudy (experimentx):

in general, i would be advisible to look for odd number of pages.

OpenStudy (shubhamsrg):

but I thought my hypothesis was a good proof that more than 3 pages is not possible :|

OpenStudy (asnaseer):

I agree with @shubhamsrg

OpenStudy (anonymous):

thanks a lot guys for all the input ,you guys are really great problem solvers ...i am out of medals...

OpenStudy (asnaseer):

@Jonask the question asks "how many pages are torn off" - so answer is 3. We cannot determine exactly which 3 but we do know that exactly 3 pages were torn off. See the reasoning outlined by @shubhamsrg above.

OpenStudy (anonymous):

yes the question is completlty answered,i understand the solution completly

ganeshie8 (ganeshie8):

can we determine how many solutions are possible

OpenStudy (asnaseer):

I think it would be the number of solutions for:\[a+b+c=49\]where a, b and c are all odd integers

ganeshie8 (ganeshie8):

yeah 3 variables so many !

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

i see number theory here i think we shud halt here

OpenStudy (asnaseer):

you can certainly solve it by brute force as the range is small :)

OpenStudy (shubhamsrg):

its a permutation and combination thing then

OpenStudy (shubhamsrg):

its surely less than 253, not sure how much less

OpenStudy (experimentx):

should be of the form a+b+c = 23 where a>0, b>0, c>0 and \( a \neq b \neq c \)

OpenStudy (shubhamsrg):

if we remove condition for a not equal b not equal c, then 25C2 or300 solutions we have but it surely has to be less than that coz of a not equals b not equals c I said 253 after assuming a>1,b>1,c>1 but that was wrong assumption, hmm..

OpenStudy (experimentx):

let a=1, b+c=22 => 19 solutions a=2, b+c= 21=> 19 solutions a=3, b+c= 20=> 18 solutions .. ..

OpenStudy (shubhamsrg):

aha nice.

OpenStudy (asnaseer):

yes - nice reduction @experimentX :)

OpenStudy (experimentx):

thanks :) still i haven't got or thought why more than three pages can't be torn off.

OpenStudy (shubhamsrg):

I'll simply copy-paste from above for n pages torn, 2(a_1+a_2+a_3 ...a_n) + n = 101 (a_1+a_2+a_3..a_n) = (101-n)/2 a_1,a_2,a_3..a_n all are odd. n also has to be odd so as to make RHS an integer. Another condition is RHS should be odd integer. n=3,7,11.. satisfy this criteria. for n=7, we have a_1 + a_2 .. a_7 = 47 sum of first 7 odd numbers is 49 i.e. greater that 47 hence anything greater than 3 should not be solution.

OpenStudy (asnaseer):

@shubhamsrg explained it nicely above...

OpenStudy (asnaseer):

and now repeated just above :)

OpenStudy (experimentx):

i see :)

OpenStudy (anonymous):

i see :) well explained ,,,